OpenStudy (anonymous):

A satelite is orbiting the Earth. In order to boost it to a higher orbit, the controllers need to increase its velocity by 200m/s(W). If the firing thrusters can provide an acceleration of 2.5m/s^2 (W), how long must they be fired in order to boost the satellite's orbit. Show full work.

OpenStudy (xishem):

Any ideas on how to approach this?

OpenStudy (anonymous):

use the knemstics equations.. its a motion question.

OpenStudy (anonymous):

*knematics

OpenStudy (xishem):

Yep. Give it a shot.

OpenStudy (anonymous):

but i tried amd didnt get it so i was wondering if u could help.

OpenStudy (xishem):

Alright, first you need to identify which equation you'll need to use. Which quantities do you know, and which are you trying to find?

OpenStudy (anonymous):

The knowns are v=200 m/s and a=2.5 and i need to find time

OpenStudy (xishem):

That's right. Now do we have any kinematics equations that use only velocity, acceleration, and time?

OpenStudy (anonymous):

yup but v2 is not given.. we need 3 knowns in order to solve but this one has only 2 knowns..so thats why i dont know what to do!

OpenStudy (xishem):

Really we aren't looking at an absolute velocity are we? We are looking at the change in velocity. We know that the change in velocity is:\[\Delta v=v_f-v_i=200\frac{m}{s}\]Does that help?

OpenStudy (anonymous):

no bc we dont have both Vs tho

OpenStudy (anonymous):

do u know how to do this?

OpenStudy (xishem):

This is the equation we're looking at, right? \[v_f=v_i+at\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

use the kinematic equations!!!!

OpenStudy (xishem):

With a little rearrangement, it becomes:\[v_f-v_i=at\]And we know that, a change in velocity is defined as:\[\Delta v = v_f-v_i\]We can plug this into our first equation to get:\[\Delta v = at \rightarrow t=\frac{\Delta v}{a}\]

OpenStudy (xishem):

When you're working with kinematics, you'll often be able to combine two quantities by some other definition. It is true that we are looking at a velocity of 200 m/s in this question. However, this is a CHANGE in velocity, not an absolute velocity. The satellite could initially be going 1000 m/s, but requires an ADDITIONAL 200 m/s to stay in orbit. Be sure to be able to differentiate between absolute velocities and relative velocities.

OpenStudy (anonymous):

SO whats the answer? I mean I know this info but i dont know how put it together and get final answer

OpenStudy (xishem):

You'll be using this equation directly:\[t=\frac{\Delta v}{a}\]which was derived from two posts above. I trust you can plug in two quantities and calculate.

OpenStudy (anonymous):

oh right! Got it! Thank you!

OpenStudy (anonymous):

80 seconds

OpenStudy (anonymous):

Can u help me with this one as well

OpenStudy (xishem):

If you've got a different question, go ahead and post it in a new question. (:

OpenStudy (anonymous):

alright thank u! :')